40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12114     Entry time: Tue May 10 03:44:59 2016
Author: ericq 
Type: Update 
Category: LSC 
Subject: Relocked 

ALSX noise is solidly within past acceptable performance levels. The DRFPMI was locked on four out of six attempts. 

Some housekeeping was done:

  • PMC aligned
  • Static alignment voltages of X end PZT mirrors offloaded by turning mount screws
  • Rough comissioning of AUX X dither alignment
  • Locking scripts reverted to AUX X Innolight voltage/temperature sign convention

The recombination of the QPD signals to common / differential is imperfect, and limited how well we could keep the interferometer aligned, since the QPD at X has changed. This needs some daytime work. 

Some sensing matrix measurements were made, to be meditated upon for how to 1F the DRMI.

Other to-dos:

  • Bandpass + notch combo for green refl PDs
  • SRCL, and to a lesser extant, MICH feedforward subtraction (See DARM vs. other length DOF coherence plot below)
  • Fiber couple AUX X light
  • Make IFO work good


As an aside, Gautam and I noticed numerous green beams coming from inside the vacuum system onto the PSL table. They exist only when green is locked to the arms. Some of them come out at very non-level angles and shine in many places. This doesn't make me feel very happy; I suppose we've been living with it for some time. 

Attachment 1: 2016-05-10_DARMcoherence.pdf  139 kB  Uploaded Tue May 10 04:51:29 2016  | Hide | Hide all
2016-05-10_DARMcoherence.pdf
ELOG V3.1.3-